OpenStudy (breee26):

Which of the following polynomials does not have real roots? A. X^2 -17 -7 B. X^2 +4x +8 C. -x^2 +4x +6 D. 2x^2 -13

OpenStudy (anthonyym):

You get a real root whenever the graph intercepts the x-axis. So you can graph each one to find out.

OpenStudy (mariamikayla):

x1+x2= -b/a If one of the root is a complex it means the sum would be a complex number,so we have for a) a=1 b= -17 => x1+x2= -(-17)/1= 17 -it has real roots b)a=1 b=4 => x1+x2= -4/1= -4 -it has real roots c)a=-1 b=4 => x1+x2= -4/-1= 4 -it has real roots d)a=2 b= 0 => x1+x2= -0/2 which doesn't make sense so the answer is variant d)

OpenStudy (anthonyym):

Another way is to use the determinant (b^2 - 4ac), from the quadratic formula. If it comes out negative there are no real roots. This works because in the quadratic formula \[x = \pm \frac{ \sqrt{b^2-4ac} }{ 2a }\], if the square root is negative it will have imaginary roots, because you can't have negative square root

OpenStudy (breee26):

Thank you @mariamikayla

OpenStudy (mariamikayla):

You're welcome :D